Which is worth more today:

  1. $450 in 10 years assuming a 4% discount rate
  1. $400 in 6 years assuming a 5% discount rate

a.Which is worth more today:

  1. $100,000 in 100 years assuming a 6% discount rate
  1. $310 in 2 years assuming a 2% discount rate

b.Which is worth more in the year given:

  1. The year is 2021: $1000 in the year 2032, assuming a 6.5% discount rate
  1. The year is 2090: $500 in the year 2108, assuming a 3% discount rate

c. In the question above (#9), what is each worth today, use the same discount rates indicated:

_________________

_________________

show all work
Answer #1

To find which is worth more today out of the two options, we need to calculate the PV of the same:

  • $450 in 10 years assuming a 4% discount rate:
    • We are given the following information
    • r 4.00%
      n 10
      frequency 1
      FV $                  450.00
    • We need to solve the following equation to arrive at the required PV:
    • FV = PV x (1+r) 450 = PV X (1+0.04)10 PV = 304.00
    • So the PV was 304.00
  • $400 in 6 years assuming a 5% discount rate
    • We are given the following information:
    • r 5.00%
      n 6
      frequency 1
      FV $                  400.00
    • We need to solve the following equation to arrive at the required FV
    • FV = PV x (1+r) 400 = PV (1+0.05) PV = 298.47
    • So the PV is 298.47
  • Therefore the first option is worth more now